OpenStudy (anonymous):

evaluate the indefinite integral (ln(x))^5/X *dx

OpenStudy (anonymous):

(Ln(x))^6 /6

OpenStudy (anonymous):

\[\int\limits \frac{\text{Log}[x]^5}{x} \, dx=\frac{\text{Log}[x]^6}{6}+c \]